OpenStudy (anonymous):

a plane starts from rest and accelerates in a straight line along the ground before takeoff. It moves 600m in 12s. Find its acceleration.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

hi i know this is easy but i just kind of got a question about finding acceleration

OpenStudy (anonymous):

so i know a=change in velocity/time

OpenStudy (amistre64):

v = at

OpenStudy (anonymous):

so the velocity would be 600/12 correct?

OpenStudy (amistre64):

m/s = m/s^2 * s

OpenStudy (amistre64):

seems fair to me

OpenStudy (anonymous):

so i dont really understand why it wouldnt just be 50/12 for acceleration

OpenStudy (anonymous):

but then if u do 50/12 then multiply that answer by 2 u get the right answer?

OpenStudy (amistre64):

spose we have a constant accleration, in order to move we need to accelerate position is: 1/2 at^2 + vo t + so velocity is at + vo acceleration is: a

OpenStudy (amistre64):

lets use position to determine acceleration since that is what we are given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are u saying vo is initial or final velocity?

OpenStudy (amistre64):

600 = 1/2 a (12)^2 1200 = a (12)^2 a = 100/12

OpenStudy (amistre64):

initial

OpenStudy (amistre64):

and so is initial position, assumed to me 0 as well

OpenStudy (anonymous):

that would be zero correct?

OpenStudy (amistre64):

yes vo and so are 0 in this case

OpenStudy (anonymous):

ok there are two more parts to this question, mind helping me out with them? u have been super helpful :D

OpenStudy (amistre64):

sure ... if my memory holds up :)

OpenStudy (anonymous):

ok sweet, the next part says find the speed at the end of 12seconds

OpenStudy (anonymous):

any idea? :D

OpenStudy (amistre64):

speed is just the velocity at t=12

OpenStudy (amistre64):

since we know acceleration: a v = at

OpenStudy (anonymous):

so its just 8.3X12?

OpenStudy (amistre64):

recall we are given a distance and time d = 1/2 a t^2 600 = 1/2 a(12)^2 100/12 = 25/3 = a

OpenStudy (amistre64):

v = 25(12)/3

OpenStudy (anonymous):

why would u divide by 3 :o

OpenStudy (anonymous):

i feel stupid for asking this stuff lol

OpenStudy (amistre64):

100/12 reduces to 25/3

OpenStudy (amistre64):

25*5/3*4 = 100/12

OpenStudy (amistre64):

a = 25/3 and t=12

OpenStudy (anonymous):

oh yea i thought 3 was like some other value or something

OpenStudy (anonymous):

and the last part said find the distance moved during the 12th second

OpenStudy (amistre64):

during the 12th second ... is that from t-12 to t=13?

OpenStudy (amistre64):

*t=12

OpenStudy (amistre64):

t=0 to 1 is the first second so t=11 to 12 is the 12th second

OpenStudy (anonymous):

yes i believe thats what it means but it just says the 12th second

OpenStudy (amistre64):

600 - 1/2 a(11)^2

OpenStudy (anonymous):

actually i think it might mean like 12th second to 13th second

OpenStudy (amistre64):

your call, im going with 11 to 12 :)

OpenStudy (amistre64):

0to1 =1 1to2 = 2 2to3 = 3 ... 11to12 = 12

OpenStudy (anonymous):

i have the answer but i need to figure out how to get to it so i guess we can just try one way

OpenStudy (anonymous):

and then the other if it doesnt work

OpenStudy (amistre64):

trying is the best way yes :)

OpenStudy (anonymous):

so what u were doing was finding acceleration for the 11th second?

OpenStudy (amistre64):

no finding the distance from t=11 to t=12 which is utilising the distance setup 1/2 a (12)^2 - 1/2 a (11)^2 = 1/2 a (12^2 - 11^2)

OpenStudy (anonymous):

so u just use that equation u made?

OpenStudy (amistre64):

lol, i didnt just 'make' it, but yes :)

OpenStudy (anonymous):

i mean like the equation u just showed me how to get lol

OpenStudy (amistre64):

if you know the distance at t=11 and the distance at t=12 ... then the difference between them is how far you hav emoved right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then lets go with 95.833333

OpenStudy (anonymous):

ok sweet man !! u helped me so much!!!!

OpenStudy (amistre64):

goodluck and all ;)